OpenStudy (cobaltthunder267):

Find sin θ, if cos θ = 2/5 and tan θ < 0

OpenStudy (shamim):

Given Cos theta=2/5

OpenStudy (shamim):

Need to draw a right angle triangle

OpenStudy (shamim):

Base=2 Hypotenuse=5

OpenStudy (shamim):

Pythagorean theorem Hypotenuse^2=base^2+perpendicular^2

OpenStudy (shamim):

5^2=2^2+perpendicular^2 Perpendicular=root 21 Right?

OpenStudy (dtan5457):

How I would go about a problem like this is... knowing that sin(theta)^2+cos(theta)^2=1 subsititute cos (2/5) sin(theta)^2+(2/5)^2=1 (2/5)^2=4/25 sin^2=21/25 take square root sin=sqrt21/5 It says that tan<0..so it's not quadrant 1 or 3, and since cos is positive, this must be the 4th quadrant and sin is negative.

OpenStudy (dtan5457):

but if you are not familiar with identities, follow @shamim 's way

OpenStudy (cobaltthunder267):

Oh, okay! I think I get it now!! I'm still working on Identities, so I wasn't quite sure how to apply them. Thank you guys so much! @dtan5457 & @shamim!

OpenStudy (dtan5457):

Alright, it's best to get familiar with both...some teachers (my teacher) forces us to use identities when asked upon...so...try and memorize them.

OpenStudy (shamim):

Welcome!!!!!